# Dark matter halo as a source of regular black-hole geometries

### v1

@@ VIII Conclusions
In this work we have constructed exact black-hole solutions free from curvature singularities, sourced by a galactic dark-matter halo. The regularity of these geometries follows from the specific relation $P_{r}=-\rho$ between radial pressure and density, which is consistent with the phenomenological freedom in halo modeling. The regular black-hole solution is shown to be stable against axial perturbations. While this could be an indication that the solution is stable, the full analysis of stability must also include the polar type of perturbations. The obtained solutions demonstrate that halo distributions can serve not only as astrophysical environments but also as effective sources of singularity resolution. The radius of shadow and the Lyapunov exponent for circular null geodesics are discussed for particular examples of regular black holes.
A natural direction for further study is the investigation of physical observables associated with these regular configurations. In particular, the analysis of gravitational spectra through quasinormal modes and echoes can provide insights into the stability and response of such black holes under perturbations. Similarly, the study of electromagnetic spectra, including gravitational lensing, and energy emission, offers potential observational signatures that could distinguish regular black holes embedded in halos from their singular Schwarzschild or Kerr counterparts.
The literature on singular black holes surrounded by galactic halos is already extensive, covering quasinormal ringing Konoplya (2021); Konoplya and Zhidenko (2022); Dubinsky (2025); Feng and Zhang (2025); Pezzella et al. (2025); Chakraborty et al. (2025); Liu et al. (2025, 2024); Zhao et al. (2023); Daghigh and Kunstatter (2022); Zhang et al. (2021), grey-body factors Hamil et al. (2025); Mollicone and Destounis (2025); Tovar et al. (2025), black-hole shadows and lensing Konoplya et al. (2025c); Hou et al. (2018); Kouniatalis et al. (2025); Fernandes and Cardoso (2025); Chen et al. (2024); Tan et al. (2025); Macedo et al. (2024); Xavier et al. (2023); Figueiredo et al. (2023); Konoplya et al. (2025d), and dynamical processes such as accretion Chowdhury et al. (2025); Heydari-Fard et al. (2025). Extending these analyses to the case of regular black holes supported by halo matter would allow one to quantify the extent to which regularity alters these spectra and potentially leaves imprints observable with current or near-future instruments.

### v3

@@ VII Discussions
While the above particular models of dark matter distributions might be questioned at the stage of black-hole formation, we do not rely on these specific choices. In fact, the emergence of a regular black-hole geometry follows from rather generic assumptions about the effective matter distribution, as outlined in Sec. II, namely, that the density remains finite and regular everywhere, and that the anisotropic fluid satisfies the weak energy condition. Unlike the vacuum case, these minimal conditions admit regular black-hole solutions for sufficiently dense matter configurations.
@@ VIII Conclusions
In this work we have constructed exact black hole solutions free from curvature singularities, sourced by a galactic dark-matter halo. The regularity of these geometries follows from the specific relation $P_{r}=-\rho$ between radial pressure and density, which is consistent with the phenomenological freedom in halo modeling. The regular black hole solution is shown to be stable against axial perturbations. While this could be an indication that the solution is stable, the full analysis of stability must also include the polar type of perturbations. The obtained solutions demonstrate that halo distributions can serve not only as astrophysical environments but also as effective sources of singularity resolution. The radius of shadow and the Lyapunov exponent for circular null geodesics are discussed for particular examples of regular black holes.
A natural direction for further study is the investigation of physical observables associated with these regular configurations Vagnozzi et al. (2023). In particular, the analysis of gravitational spectra through quasinormal modes and echoes can provide insights into the stability and response of such black holes under perturbations. Similarly, the study of electromagnetic spectra, including gravitational lensing Boos and Hu (2025), and energy emission, offers potential observational signatures that could distinguish regular black holes embedded in halos from their singular Schwarzschild or Kerr counterparts.
The literature on singular black holes surrounded by galactic halos is already extensive, covering quasinormal ringing Konoplya (2021); Konoplya and Zhidenko (2022); Dubinsky (2025); Feng and Zhang (2025); Pezzella et al. (2025); Chakraborty et al. (2025); Liu et al. (2025, 2024); Zhao et al. (2023); Daghigh and Kunstatter (2022); Zhang et al. (2021), grey-body factors Hamil et al. (2025); Mollicone and Destounis (2025); Tovar et al. (2025); Lütfüoğlu (2025b); Pathrikar (2025), black hole shadows and lensing Konoplya et al. (2025c); Hou et al. (2018); Kouniatalis et al. (2025); Fernandes and Cardoso (2025); Chen et al. (2024); Tan et al. (2025); Macedo et al. (2024); Xavier et al. (2023); Figueiredo et al. (2023); Konoplya et al. (2025d), and dynamical processes such as accretion Chowdhury et al. (2025); Heydari-Fard et al. (2025). Extending these analyses to the case of regular black holes supported by halo matter would allow one to quantify the extent to which regularity alters these spectra and potentially leaves imprints observable with current or near-future instruments.